Marlow Company purchased a point of sale system on January 1 for $3,400. This system has a useful life of 10 years and a salvage value of $400. What would be the depreciation expense for the second year of its useful life using the double declining balance method?
Answer #1

Calculation of annual depreciation rate:

Annual depreciation rate= 100%/number of years= 100%/10= 10%

Double declining rate= 10*2= 20%

Depreciation for year 1= 3400*0.20= 680

Depreciation for year 2= (3400-680)*0.20= 544

Depreciation expense for the second year is $544